A body is projected from the top of a tower with a velocityu=3i+4j+5km/s,where i,j,k are unit vectors along east,northand vertically upwards respectively.If the height of the tower is 30m, horizontal range of the bodyon the ground is
 
 

Grade:9

2 Answers

Aman Kashyap
6 years ago
We need to find range so we dont need north and east direction se we can write the magnitude of 3i+4j as 5m/s. And we have given 5 m/s upward(k). We are also given highy of the tower so we can easily calculate the time of flight. First we apply 2u/g formula for the time, when the particle goes up and then comes down at the hight of the tower. 2(5)/10 = 1second. Now next step is to calulate the speed at this moment, which can be orally told to be 5m/s because it was thrown up with 5m/s. Now we apply S = ut + (1/2)at^2 and put 30m instead of S, 5 instead of u and 10 instead of a. It will form a quadratic in t : t^2 + t - 30 = 0. t=5. Now total time is 6 sec. So range =6×5=30m. Thanks Approve if it helped and good luck for you future
dmee
2 years ago

A ball is projected From the top of a tower with a. velocity 3i+ 4j- 5k m/s,where i,j,k are unit vectors along east,north and vertical upwards respectively.If the height of the tower is 30m,its range is?